OpenStudy (anonymous):

Here is the graph of an exponential function corresponding to a geometric sequence.

OpenStudy (anonymous):

Which of the following is the explicit formula for the geometric sequence? A. an = 9(one third)n+1 B. an = 9(one third)n-1 C. an = 9(3)n-1 D. an = 9(3)n+1

OpenStudy (anonymous):

@xapproachesinfinity